# CERN-TH-2024-177

### v1

@@ Appendix B Particle Swarm implementation
To perform the search we first choose a region in the fit parameter space where we expect to find the minimum of the $\chi^{2}(\Theta)$ function. In our case, we choose a region located around the phenomenological estimates of the fit parameters. The position of each particle is initialized with random numbers drawn from a Gaussian distribution, with the mean at the center of the region and variance proportional to its size, which we denote by $\vec{\epsilon}$. As for the number of particles, we use $n_{\text{p}}=10$. We do a similar operation for the initial velocities of the particles: we pick them from a Gaussian distribution with zero mean and variance $\vec{\epsilon}$.
Then we start the algorithm by updating the position and velocities of the particles according to (11). We also enforce the rule that when a particle is sent outside of the box by the iteration step, we place it back on the wall with the opposite velocity.
To efficiently minimize the $\chi^{2}(\Theta)$ we use a small variation of the standard PSO called “adaptive PSO” which uses the velocity information to tune the value of $\omega$, as described in Xu (2013). We use $c_{1}=c_{2}=1.4962$, and vary $\omega$ ranging from $0.3\leq\omega\leq 0.8$ starting from the highest value $0.8$, and then decrease it or increase it at each step depending on the value of the average velocity of the swarm.
Without the adaptive step, the swarm will run until it reaches an agreement on the best position found and eventually will start exploring the neighbor of the optimal one. Since, in our case, evaluating $\chi^{2}(\Theta)$ is as computationally expensive as an entire S-matrix Bootstrap run, we adopt a strategy to reduce the number of steps required for convergence.
The idea is to tune the friction parameter $\omega$ according to some control function. The friction $\omega$ controls how well the swarm will explore the parameter space. If $\omega$ were of order one, the particles would wildly explore the space with the risk of never agreeing on the minimum. If $\omega$ were too small, the particles would immediately freeze without exploring much.
In Xu (2013), it is suggested to look at the total velocity of the swarm at each step $n$ defined as
@@
If $\mathbb{V}_{n-1}\geq\mathbb{V}^{\text{ideal}}_{n}$, we set $\omega_{n}=\max(\omega_{n-1}-\Delta\omega,\omega_{\text{min}})$, otherwise $\omega_{n}=\min(\omega_{n-1}+\Delta\omega,\omega_{\text{max}})$. In our case, $\omega_{\text{max}}=0.8$, $\omega_{\text{min}}=0.3$, and $\Delta\omega=0.1$.
@@
The danger of this approach is that if we choose $N_{\text{iter}}$ too small, we might end up in a local minimum. For this reason, we performed two searches with different $N_{\text{iter}}=30,60$.
In figure 13 we plot the positions of the particles in the $(z_{0},z_{2})$ plane at different steps $n$ in the run with $N_{\text{iter}}=60$. The black dots correspond to the actual positions of the particles at step $n$. The blue dots correspond to each particle’s best position up to step $n$, and the red dot is the best found by the swarm at the same step. At $n=0$, black and blue dots coincide. The gray area is the box where we perform the search.

### v2

@@ Appendix B Particle Swarm implementation
To begin the algorithm, we first choose a search region in the fit parameter space where we will look for the minimum of the $\chi^{2}(\Theta)$ function. Specifically, we select a region $\vec{\Theta}_{\text{min}}\leq\vec{\Theta}\leq\vec{\Theta}_{\text{max}}$ centered around the phenomenological estimates of the fit parameters.
We set the number of particles to $n_{\text{p}}=10$. This number represents the size of the “swarm”. The initial positions of the particles are generated randomly, with values drawn from a Gaussian distribution centered at the region’s center, and a variance proportional to its size, denoted by $\vec{\epsilon}=\tfrac{1}{2}(\vec{\Theta}_{\text{max}}-\vec{\Theta}_{\text{min}})$. Similarly, the initial velocities of the particles are sampled from a Gaussian distribution with zero mean and variance $\vec{\epsilon}$.
We start the algorithm by updating the positions and velocities of the particles according to equation (11). Additionally, if the particle $i$ moves outside the boundaries of the box during the update step along some direction $j$, we then place it back on the boundary setting $(\Theta_{n+1}^{(i)})_{j}=(\Theta^{(i)}_{\text{max/min}})_{j}$ with the opposite velocity.
Evaluating $\chi^{2}(\Theta)$ for a given set of fit parameters is computationally expensive, as it involves solving an S-matrix bootstrap problem. To efficiently estimate the minimum of the $\chi^{2}(\Theta)$ function in fewer steps, we use a modified version of the standard PSO, called adaptive PSO. We use velocity information to adjust the value of $\omega$, as described in Xu (2013).
The goal is to adjust the friction parameter $\omega$ using a control function. The value of $\omega$ determines how effectively the swarm explores the parameter space. If $\omega$ is too large (of order one), the particles will explore erratically, risking failure to converge on the minimum. If $\omega$ is too small, the particles will freeze early, preventing a broad exploration.
In Xu (2013), the authors suggest to look at the total velocity of the entire swarm at each step $n$ defined as
@@
If $\mathbb{V}_{n-1}\geq\mathbb{V}^{\text{ideal}}_{n}$, we set $\omega_{n}=\max(\omega_{n-1}-\Delta\omega,\omega_{\text{min}})$, otherwise $\omega_{n}=\min(\omega_{n-1}+\Delta\omega,\omega_{\text{max}})$. In our case, $\omega_{\text{max}}=0.8$, $\omega_{\text{min}}=0.3$, and $\Delta\omega=0.1$. We also set $c_{1}=c_{2}=1.5$.
@@
The risk with this approach is that choosing a small $N_{\text{iter}}$ could lead to getting stuck in a local minimum. To mitigate this, we performed two searches with different values of $N_{\text{iter}}=30$ and 60.
Figure 14 shows the projection of particle positions onto the $(z_{0},z_{2})$ plane at various steps $n$ during a run with $N_{\text{iter}}=60$. Black dots represent the particles’ current positions $\Theta_{n}^{(i)}$ at step $n$, blue dots indicate each particle’s best position $X_{n}^{(i)}$ up to that step, and the red dot marks the global best position $Y_{n}$ found by the swarm. At $n=0$, the black and blue dots coincide. The gray region represents the search domain.